A motorboat can maintain a constant speed of 35 miles per hour relative to the water. The boat makes a trip upstream to a certain point in 48 minutes; the return trip takes 32 minutes. What is the speed of the current?
Answer by CubeyThePenguin(3113)

speed of current = C

distance = (rate)(time)

(35 + C)(32/60) = (35 - C)(48/60)
8(35 + C)/15 = 4(35 - C)/5
280 + 8C = 3(140 - 4C)
280 + 8C = 420 - 12C
20C = 140
C = 7

The current travels at 7 mph.